HIGH RESOLUTION INFRARED SPECTRA OF PLASMA JET-COOLED DI- AND TRIACETYLENE IN THE C-H STRETCH REGION BY CW CAVITY RING-DOWN SPECTROSCOPY

\underline{D. ZHAO}, J. GUSS, A. WALSH, K. DONEY, H. LINNARTZ, Sackler Laboratory for Astrophysics, Leiden Observatory, University of Leiden, P.O. Box 9513, NL-2300 RA Leiden, the Netherlands.

Polyacetylenes form an important series of unsaturated hydrocarbons that are of astrophysical interest. Small polyacetylenes have been detected from infrared observations in dense atmosphere of Titan and in a protoplanetary nebula CRL 618. We present here high-resolution mid-infrared spectra of diacetylene (HC_{4}H) and triacetylene (HC_{6}H) that are recorded in a supersonically expanded pulsed planar plasma using an ultra-sensitive detection technique. This method uses an all fiber-laser-based optical parametric oscillator (OPO), in combination with continuous wave cavity ring-down spectroscopy (cw-CRDS) as a direct absorption detection tool^{a}. A hardware-based multi-trigger concept is developed to apply cw-CRDS to pulsed plasmas. Vibrationally hot but rotationally cold HC_{4}H and HC_{6}H are produced by discharging a C_{2}H_{2}/He/Ar gas mixture which is supersonically expanded into a vacuum chamber through a slit discharge nozzle. Experimental spectra are recorded at a resolution of ∼100 MHz in the 3305-3340 cm^{−1} region, which is characteristic of the C-H stretch vibrations of HC_{4}H and HC_{6}H. Jet-cooling in our experiment reduces the rotational temperature of both HC_{4}H and HC_{6}H to <20 K. In total, ∼2000 lines are measured. More than fourteen (vibrationally hot) bands for HC_{4}H and four bands for HC_{6}H are assigned based on Loomis-Wood diagrams, and nearly half of these bands are analyzed for the first time. For both molecules improved and new molecular constants of a series of vibrational levels are presented. The accurate molecular data reported here, particularly those for low-lying (bending) vibrational levels may be used to interpret the ro-vibrational transitions in the FIR and submillimeter/THz region.

aD. Zhao, J. Guss, A. Walsh, H. Linnartz Chem. Phys. Lett., \underline{dx.doi.org/10.1016/j.cplett.2013.02.025}, in press, 2013.